### v4

# Xinlu Zhang1, Chenxin Tian2, Xianjun Yang1, Lichang Chen3, Zekun Li1, Linda Ruth Petzold 1

Abstract
Instruction-finetuning (IFT) has become crucial in aligning Large Language Models (LLMs) with diverse human needs and has shown great potential in medical applications. However, previous studies mainly fine-tune LLMs on biomedical datasets with limited diversity, which often rely on benchmarks or narrow task scopes, and hence significantly limit the effectiveness on their medical instruction-following ability and generalizability. To bridge this gap, we propose creating a diverse, machine-generated medical IFT dataset, MedInstruct-52k, using GPT-4 and ChatGPT with a high-quality expert-curated seed set. We then fine-tune LLaMA-series models on the dataset to develop AlpaCare. Despite using a smaller domain-specific dataset than previous medical LLMs, AlpaCare not only demonstrates superior performance on medical applications, with up to 38.1% absolute gain over best baselines in medical free-form instruction evaluations, but also achieves 6.7% absolute gains averaged over multiple general domain benchmarks. Human evaluation further shows that AlpaCare consistently outperforms best baselines in terms of both correctness and helpfulness. We offer public access to our data, model, and codebase in https://github.com/XZhang97666/AlpaCare.
1 Introduction
Recent advancements in the training of large language models (LLMs) have placed a significant emphasis on instruction-finetuning (IFT), a critical step in enabling pre-trained LLMs to effectively follow instructions (Ouyang et al., 2022; Longpre et al., 2023; Wei et al., 2022). However, relying solely on NLP benchmarks to create instructional datasets can lead to ‘game-the-metric’ issues, often failing to meet actual user needs (Ouyang et al., 2022). To better align with human intent, Wang et al. (2023b) introduces the concept of fine-tuning LLMs using diverse machine-generated instruction-response pairs. Subsequent works further highlight the importance of diversity in IFT datasets (Taori et al., 2023; Xu et al., 2023a; Chiang et al., 2023; Chen et al., 2023; Bukharin & Zhao, 2024). However, how to improve dataset diversity in the medical domain for aligning with various user inquiries is still underexplored.
LLMs have demonstrated significant potential in the medical domain across various applications (Nori et al., 2023a; b; Singhal et al., 2022; 2023; Liévin et al., 2023; Zhang et al., 2023). To alleviate privacy concerns and manage costs, several medical open-source LLMs (Han et al., 2023; Li et al., 2023b; Wu et al., 2023; Xu et al., 2023b) have been developed by tuning LLaMA (Touvron et al., 2023a; b) on medical datasets. Even substantial volumes, these datasets are limited in task scopes and instructions, primarily focusing on medical benchmarks or specific topics, due to the high cost of collecting real-world instruction datasets (Wang et al., 2023b), particularly when extending further into the medical domain(Jin et al., 2021; 2019). This lack of diversity can negatively impact the models’ ability to follow instructions in various medical applications and their effectiveness in the general domain. Therefore, there is an urgent need for a method to generate diverse medical IFT datasets that align with various domain-specific user inquiries while balancing cost.
To bridge this gap, inspired by Wang et al. (2023b), we propose a semi-automated process that uses GPT-4 (OpenAI, 2023) and ChatGPT (OpenAI, 2022) to create a diverse medical IFT dataset for tuning a medical LLM, which can better align with various domain-specific user intents. Initially, to guide the overall task generations with meaningful medical instructions and considering different user needs, we create a high-quality seed set of 167 clinician-curated tasks spanning various medical topics, points of view, task types, and difficulty levels, as shown in Figure 1.
Figure 1: Selected example from the clinician-crafted seed set. We focus on 4 perspectives: topic, viewpoint, task type, and difficulty level, to improve the seed set diversity. The set is further used to query GPT-4 to generate medical tasks.
To automatically generate a broader array of tasks for training, we prompt GPT-4 to create instructions for new medical tasks by leveraging the existing clinician-curated tasks as demonstrations. After generating tasks and conducting deduplications, we employ ChatGPT to provide responses to the valid tasks. Consequently, we compile a 52k medical self-instruct dataset, MedInstruct-52k, which supervises the tuning on the LLaMA series models (Touvron et al., 2023a; b), resulting in AlpaCare. Due to the limited number of test sets available for evaluating medical LLMs in terms of instruction-following ability and medical capacity, we introduce a new clinician-crafted free-form instruction evaluation test set, MedInstruct-test, covering medical tasks across different difficulty levels.
Our comprehensive experiments within medical and general domains reveal that AlpaCare, solely tuned on the 52k diverse medical IFT dataset, exhibits enhanced performance on medical applications and strong generalizability. It achieves up to a 38.1% absolute gain over the best baselines in medical free-form instruction evaluations and a 6.7% absolute gain averaged over multiple general domain benchmarks. Moreover, our human study on free-form instruction evaluations shows that AlpaCare consistently produces better responses compared to existing medical LLMs by a large margin in terms of both correctness (+12%) and helpfulness (+49%).
Our paper makes the following contributions:
  • To address the challenge of generating cost-effective, high-quality IFT data for LLM alignment in various medical applications, we propose a pipeline to create a diverse medical machine-generated IFT dataset for tuning LLMs.
  • We conduct extensive experiments in medical and general domains, demonstrating that tuning LLMs with a diverse medical IFT dataset can boost their capacity in medical applications and generalization simultaneously.
  • We release, MedInstruct-52K, a diverse medical IFT dataset comprising 52K instruction-response pairs and, MedInstruct-test, a test set of 216 clinician-crafted novel medical tasks, to facilitate the building and evaluation of medical LLMs.
2 Related Works
IFT. Closed-form IFT (Wei et al., 2022; Longpre et al., 2023) creates IFT datasets from existing NLP benchmarks to improve model generalization on new tasks. For instance, FLAN (Wei et al., 2022), T0 (Sanh et al., 2022), and Flan-T5 (Chung et al., 2022) construct their IFT datasets using existing NLP tasks with carefully designed instructions and demonstrate that fine-tuning with diverse task instructions enhances performance on unseen tasks. However, these closed-form instructions are often simpler compared to cases in real-world scenarios, making their models fail to align with various real-world user intentions. Alternatively, Ouyang et al. (2022) collects a diverse IFT dataset with real-world instructions and responses, which is rich in both instruction forms and task types, and trains GPT-3 Brown et al. (2020) to obtain InstructGPT (Ouyang et al., 2022) on this dataset, showing promising results in aligning with diverse actual user needs. Due to the closed-source propriety of strong LLMs (e.g. ChatGPT and GPT-4), various open-source instruction fine-tuned models (Taori et al., 2023; Xu et al., 2023a; Chiang et al., 2023; Peng et al., 2023) have been proposed to fine-tune open-source LLMs using datasets obtained from these strong teacher models to enhance their instruction-following abilities . Alpaca (Taori et al., 2023) creates a 52k diverse machine-generated IFT dataset by distilling knowledge from the "teacher" Text-Davinci-003 (Hinton et al., 2015; Li et al., 2022). Peng et al. (2023) utilizes the same instructions with Alpaca but adopts GPT-4 as the "teacher" LLM to generate higher-quality and more diverse responses to improve the model’s alignment on 3H (Helpfulness, Honesty, and Harmlessness) (Askell et al., 2021). Vicuna (Chiang et al., 2023) is trained on the ShareGPT data (Sharegpt, 2023), which contains actual ChatGPT users’ diverse instructions, obtaining strong response quality and instruction-following ability. However, creating diverse IFT datasets for aligning models with various domain-specific user intentions in the medical domain remains underexplored.
LLMs in Biomedicine. Closed-source LLMs have demonstrated significant proficiency in the medical domain (OpenAI, 2022; 2023; Singhal et al., 2023; 2022). ChatGPT demonstrates promise in the US Medical Exam (Kung et al., 2023) and serves as a knowledge base for medical decision-making (Zhang et al., 2023). The MedPaLM (Singhal et al., 2022; 2023) have shown performance in answering medical questions on par with that of medical professionals. GPT-4 (OpenAI, 2023) obtains strong medical capacities without specialized training strategies on medical domain or engineering for solving clinical tasks Nori et al. (2023a; b). Due to privacy concerns and high costs, several open-source medical LLMs (Xu et al., 2023b; Han et al., 2023; Li et al., 2023b; Wu et al., 2023) have be built by tuning open-source base model, such as LLaMA (Touvron et al., 2023a; b), on medical corpus. However, due to the high cost of collecting diverse real-world user instructions (Wang et al., 2023b), their datasets are limited in diversity and primarily focus on medical benchmarks or narrow task scopes, such as doctor-patient conversations. ChatDoctor (Li et al., 2023b) is fine-tuned using 100k online doctor-patient dialogues, while Baize-Healthcare (Xu et al., 2023b) employs about 100k Quora and MedQuAD dialogues. MedAlpaca (Han et al., 2023) utilizes a 230k dataset of question-answer pairs and dialogues. PMC-LLAMA (Wu et al., 2023) continually trains LLaMA with millions of medical textbooks and papers, and then tunes it with a 202M-token dataset formed by benchmarks and dialogues in the IFT stage. However, these datasets are limited in diversity, mainly focusing on benchmarks or within certain topics, hampering models’ medical instruction-following ability and generalizability. To address this, we propose building a cost-effective diverse medical machine-generated IFT dataset by using GPT-4 and ChatGPT for model tuning to better align the model with various medical user intents. Our follow-up works (Xie et al., 2024; Tran et al., 2023) consistently show the benefits of tuning medical LLMs with diverse machine-generated datasets.
Figure 2: The pipeline of AlpaCare. The process starts with a small set of clinician-curated seed tasks. 1. Task instruction generation: GPT-4 iteratively generates a series of new task instructions using 3 tasks from the seed set. 2. Filtering: Ensures textual diversity by removing similar instructions via Rouge-L. 3. Response generation: ChatGPT creates responses for each task, forming MedInstruct-52K. 4. Instruction tuning: The dataset is used to fine-tune LLaMA models, developing AlpaCare.
3 Method
Collecting a large-scale medical IFT dataset is challenging because it necessitates 1) a deep understanding of the specific domain knowledge and 2) creativity in designing novel and diverse tasks by considering different real-world medical needs. To mitigate human effort while maintaining high quality, we propose a pipeline by instructing GPT-4 and ChatGPT to create a machine-generated dataset containing diverse domain-specific tasks. The process starts with utilizing a small set of high-quality clinician-curated seed tasks with 167 instances to prompt GPT-4 in generating medical tasks. Similar instructions are removed from the generated medical tasks, preserving 52k instances which are subsequently inputted into ChatGPT for response generation. The instruction-response pairs dataset, MedInstruct-52k, is used to tune the LLaMA, resulting in AlpaCare with superior medical instruction-following ability and generalizability. The pipeline is shown in Figure 2.
3.1 Clinician-curated seed dataset
A diverse and high-quality seed task is essential for prompting LLMs in automatic task generation (Wang et al., 2023b). We focus on 4 key areas, taking into account various user intents in medical applications, to improve the diversity of seed instructions: topic, view, type, and difficulty level. Specifically, the topic covers various submedical domains, such as radiology, genetics, and psychophysiology. The view is derived from diverse medical personnel, including researchers, medical students, and patients, who have different inquiries, to ensure a comprehensive range of viewpoints based on various levels of domain knowledge. For the type, we include various task formats, such as summarization, rewriting, single-hop, and multi-hop reasoning, to align with different application needs. Lastly, each task is categorized by its medical difficulty level, ranging from 1 to 5 (low to high), to ensure that the seed tasks can prompt new tasks on a wide range of expertise levels. We defer the explanation of the difficulty score to Appendix for further clarification. A clinician crafts each task considering these 4 dimensions, and each task contains instruction and may have a corresponding input, which could be a detailed medical example to further elucidate the instruction and enhance task diversity. Examples are shown in Figure 1.
3.2 Medical IFT dataset genreation and LLM tuning
We utilize GPT-4 for in-context learning by randomly selecting 3 tasks from the seed set and generating 12 tasks for each run. To ensure generated task diversity, we instruct GPT-4 to consider the 4 aspects outlined in 3.1. Detailed prompt instructions are provided in Table 10 in the Appendix. To further amplify textual diversity, instructions with a Rouge-L similarity above 0.7 to any other generated task are discarded (Wang et al., 2023b). Due to the lengthy propriety of medical text, we separately generate responses for each task using ChatGPT (GPT-3.5-turbo), which has demonstrated efficacy in the medical domain (Zhang et al., 2023). Finally, we result 52k machine-generated medical instruction-response pairs, MedInstruct-52k.
In IFT stage, we adopt the same training prompt and hyper-parameter setup as Taori et al. (2023) to fine-tune LLaMA models on MedInstruct-52k, Specifically, we employ instructions and inputs (when available) as inputs to tune the model to generate corresponding response outputs through a standard supervised fine-tuning with cross-entropy loss. We defer hyper-parameter setup into Appendix A.3.
4 Experimental Setup
4.1 Free-form Instruction Evaluation
Datasets. To evaluate the effectiveness of LLMs in a user-oriented manner, we conduct free-form instruction evaluations on two medical datasets. (1) iCliniq [^1] We randomly selected 1,000 instances for evaluation from the 10,000 instances proposed by Li et al. (2023b)., a dataset comprising transcripts of real patient-doctor conversations collected from an online website (Li et al., 2023b). In this task, the model processes patient inquiries as input and then simulates a doctor to provide corresponding answers. (2) MedInstruct-test, a dataset created by our clinicians, includes 216 medical instructions. These instructions mimic inquiries posed by different medical personnel, varying in difficulty on a scale from 1 to 5, with 1 being the simplest and 5 being the most challenging. We present the statistics of this dataset in Table 1, and the description of the difficulty levels is provided in the Appendix A.1.
Table 1: MedInstruct-test statistics. The distribution of task counts across various difficulty levels in MedInstruct-test is approximately equal to comprehensively evaluate medical proficiency.
Difficulty Level 1 2 3 4 5
Count 44 46 41 41 44
Evaluation Metric. We conduct auto-evaluation by employing GPT-3.5-turbo to serve as a judge (Zheng et al., 2023). The judge pairwise compares responses from a model with reference responses produced by another LLM API for each instruction in the test sets. To conduct a holistic evaluation, we employ reference outputs generated by 4 different APIs: Text-davinci-003, GPT-3.5-turbo, GPT-4 and Claude-2, respectively. To ensure unbiased evaluation and avoid positional bias (Wang et al., 2023a), we evaluate each output comparison twice, alternating the order of the model output and the reference output. We follow Li et al. (2023a) to score our models and baselines by calculating the win rate. To ensure fair comparisons, we set the maximum token length to 1024 and utilize greedy decoding for the generation of all model outputs and reference responses.
4.2 Benchmark Evaluation
Datasets. We further evaluate AlpaCare on 4 medical multiple-choice benchmarks, namely MedQA (Jin et al., 2021), HeadQA (Vilares & Gómez-Rodríguez, 2019), PubmedQA (Jin et al., 2019), and MedMCQA (Pal et al., 2022), as well as a summarization dataset, i.e., MeQSum (Ben Abacha & Demner-Fushman, 2019) [^2] We randomly selected 200 out of 1000 instances in MeQSum., to assess the model’s medical capacity.
Evaluation Metric. Following Gao et al. (2021), we conduct the multiple-choice benchmark evaluation and report the accuracy. For the summarization task, we utilize greedy decoding with a maximum token length of 1024 to generate outputs and report the ROUGE-L score.
4.3 Baselines
We evaluate the performance of AlpaCare by comparing it with both general and medical LLMs based on the LLaMA models. We consider a range of models including: (1) Alpaca, tuning on 52k general domain machine-generated samples with responses from Text-davinci-003; (2) ChatDoctor, fine-tuning with 100k real patient-doctor dialogues; (3) MedAlpaca, utilizing approximately 230k medical instances such as Q&A pairs and doctor-patient conversations; (4) PMC-LLaMA (PMC), a two-step tuning model that was first trained on 4.8 million biomedical papers and 30k medical textbooks, then instruction-tuned on a corpus of 202 million tokens; and (5) Baize-Healthcare (Baize-H), training with around 100k multi-turn medical dialogues.
Table 2: Comparative analysis of free-form instruction evaluation. Performance comparison of AlpaCare and instruction-tuned baselines. GPT-3.5-turbo acts as a judge for pairwise auto-evaluation. Each instruction-tuned model is compared with 4 distinct reference models: Text-davinci-003, GPT-3.5-turbo, GPT-4, and Claude-2. ‘AVG’ denotes the average performance score across all referenced models in each test set.
iCliniq MedInstruct
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
Alpaca 38.8 30.4 12.8 15.6 24.4 25.0 20.6 21.5 15.6 22.5
ChatDoctor 25.4 16.7 6.5 9.3 14.5 35.6 18.3 20.4 13.4 18.2
Medalpaca 35.6 24.3 10.1 13.2 20.8 45.1 33.5 34.0 29.2 28.1
PMC 8.3 7.2 6.5 0.2 5.5 5.1 4.5 4.6 0.2 4.6
Baize-H 41.8 36.3 19.2 20.6 29.5 35.1 22.2 22.2 15.6 26.6
AlpaCare 66.6 50.6 47.4 49.7 53.6 67.6 49.8 48.1 48.4 53.5
5 Experiment Results
5.1 Main Results
Free-form Instruction Evaluation Performance. The evaluation results for 4 reference models on both datasets are summarized in Table 2. AlpaCare outperforms its general domain counterpart, Alpaca, demonstrating that domain-specific training bolsters medical capabilities. Despite tuning with only 52k medical instruction-response pairs, AlpaCare consistently and significantly surpasses other medical models, which are trained on considerably larger datasets, across various reference LLMs. Specifically, for average scores across reference models, AlpaCare demonstrates a relative gain of 130% on iCliniq and 90% on MedInstruct, respectively, compared to the best baselines. These results highlight the advantages of improving medical proficiency by training with a diverse, domain-specific IFT dataset. Surprisingly, medical LLMs don’t always outperform general ones in medical tasks, and some even fail to generate useful responses, possibly due to their limited training scope restricting conversational skills.
Table 3: Results on medical benchmarks. ‘AVG’ represents the mean performance score across tasks.
MEDQA HeadQA PubmedQA MEDMCQA MeQSum AVG
Alpaca 35.7 29.1 75.4 29.2 24.4 38.8
ChatDoctor 34.3 30.0 73.6 33.5 27.1 39.7
Medalpaca 38.4 30.3 72.8 31.3 11.0 36.8
PMC 34.2 28.1 68.2 26.1 9.2 33.2
Baize-H 34.5 29.3 73.8 32.5 8.1 35.6
AlpaCare 35.5 30.4 74.8 33.5 29.0 40.6
Benchmark Evaluation Performance. Table 3 presents an extensive evaluation of AlpaCare on 5 medical benchmarks. AlpaCare obtain the best performance on average, highlighting its robust capability in the medical domain. Benchmarks evaluate a model’s intrinsic knowledge(Gao et al., 2021), which is mainly gained in LLM pretraining instead of instruction fine-tuning (Zhou et al., 2023). AlpaCare’s strong medical capability, combined with its superior ability to follow medical instructions, enables it to meet a wide range of medical application needs effectively.
5.2 Generalizability Evaluation
Training models with specific data may lead to catastrophic forgetting, limiting their generalizability (Kirkpatrick et al., 2017). Our approach, instruction tuning a model with a diverse, domain-specific dataset, aims to improve its generalizability simultaneously. We test this using AlpaCare in AlpaFarm (Dubois et al., 2023), MMLU (Hendrycks et al., 2021), BBH (Suzgun et al., 2022) and TruthfulQA (Lin et al., 2022). We compare AlpaCare with 4 reference LLMs in AlpaFarm and report the average score. The detailed score for each reference model is deferred to Table 13 in the Appendix. We follow Chia et al. (2023) to holistically evaluate models’ general domain knowledge on MMLU (5-shot) and BBH (3-shot), receptively; and evaluate the model truthfulness on TruthfulQA (0-shot) with Gao et al. (2021). The results are shown in Table 4.
Table 4: Performance on general domain tasks. AlpacaFarm is a free-form instruction evaluation, MMLU and BBH are knowledge benchmarks and TruthfulQA is a truthfulness task. ‘AVG’ denotes the average score across all tasks.
AlpacaFarm MMLU BBH TruthfulQA AVG
Alpaca 22.7 40.8 32.4 25.6 30.4
ChatDoctor 21.2 34.3 31.9 27.8 28.8
Medalpaca 25.8 41.7 30.6 24.6 30.7
PMC 8.3 23.6 30.8 23.8 21.6
Baize-H 18.3 36.5 30.1 23.5 27.1
AlpaCare 40.7 45.6 34.0 27.5 37.0
Medical LLMs typically lag behind or achieve results comparable to the general domain LLM, Alpaca, in terms of generalizability. However, AlpaCare significantly outperforms both medical and general domain baselines in multiple general tasks on average. Specifically, AlpaCare shows a significant relative improvement of 57.8% on AlpacaFarm compared to the best baseline, demonstrating strong general instruction-following ability. Moreover, AlpaCare scores higher in general knowledge tasks and maintains comparable truthfulness scores compared to other baselines, indicating strong generalization abilities due to high data diversity.
5.3 Ablation Study
To further understand the effectiveness of AlpaCare, we conduct systematic ablation studies on two medical free-form instruction evaluations and report the average score of each task across 4 reference models, receptively. The results of each reference model are detailed in the Appendix A.4.2.
AlpaCare consistently delivers superior performance in 13B model comparisons. To explore the impact of scaling up the LLM backbone, we fine-tune AlpaCare-13B on LLaMA-13B and compare its performance against other 13B LLMs.The results are shown in Table 6.
AlpaCare-13B consistently outperforms other 13B models in both tasks. This reaffirms the conclusion drawn from the 7B model comparison: tuning models with a diverse medical IFT dataset can better align the model with user needs across different medical applications.
Table 5: Result comparison on 13B instruction-tuned models. Comparing the average performance of AlpaCare-13B and its 13B baselines with GPT-3.5-turbo as the judge across 4 distinct reference models.
Alpaca Medalpaca PMC AlpaCare
iCliniq 31.3 3.9 25.4 54.4
MedInstruct 26.9 0.1 34.7 54.5
Table 6: Results on different LLM backbones. Comparing the performance of Alpaca and AlpaCare using different LLM backbones.
iCliniq MedInstruct
LLaMA Alpaca 24.4 23.2
AlpaCare 53.6 53.5
LLaMA-2 Alpaca 30.3 26.8
AlpaCare 53.7 54.2
AlpaCare achieves superior performance across various backbones. To explore the effect of different LLM backbones, we tune Alpaca-LLaMA2 and AlpaCare-LLaMA2 by training LLaMA2-7B on Alpaca data and MedInstruct-52k, respectively. Table 6 compares the performance of Alpaca and AlpaCare based on different LLM backbone families.
Consistent with the results of using LLaMA-1 as the backbone, AlpaCare-LLaMA2 consistently and significantly outperforms Alpaca-LLaMA2 in both datasets. This further underscores the backbone agnostic property of our method and emphasises tuning with a diverse medical IFT dataset can bolsters models’ medical capabilities.
Table 7: Results evaluated by the different judge. Free-form instruction evaluation with Claude-2 as the judge.
iCliniq MedInstruct
Alpaca 26.7 23.5
ChatDoctor 17.4 21.7
Medalpaca 26.7 23.1
PMC 1.3 1.8
Baize-H 25.5 19.8
AlpaCare 38.8 31.5
AlpaCare shows robust performance across different judges. Recent studies have highlighted potential biases in the LLM evaluator (Wang et al., 2023a). ChatGPT may give a higher preference for outputs from ChatGPT and GPT-4, which are both trained by OpenAI. To robustly evaluate our method, we introduce an alternative judge, Claude-2 (Anthropic, 2023) from Anthropic, to mitigate the potential biases of relying on a single family of judges. The results are shown in Table 7.
Upon evaluation by Claude-2, it is observed that AlpaCare consistently outperforms its IFT baselines by a large margin. This aligns with findings from assessments using GPT-3.5-turbo as the judge. Such consistency underscores the superior medical proficiency of our approach.
6 Human Study
Figure 3: Human study results. Head-to-head clinician preference comparison between AlpaCare-13B and PMC-13B on 100 medical free-form instruction instances on (a) correctness and (b) helpfulness.
We further conduct human studies to label question-and-answer pairs in medical free-form instruction evaluation. Three annotators with MD degrees in progress are involved in the study to perform pairwise comparisons for each question and answer pair. Specifically, we randomly select 50 prompts from each test set, totaling 100 prompts. These prompts, along with the responses generated by both AlpaCare-13B and PMC-13B, the best baseline in the 13B models, are presented to the annotators for evaluation. The evaluation is based on two criteria: correctness and helpfulness. Correctness evaluates whether the response provides accurate medical knowledge to address the question posed, while helpfulness measures the model’s ability to assist users concisely and efficiently, considering the user intent. In practical terms, an answer can be correct but not necessarily helpful if it is too verbose and lacks guidance. To determine the final result for each criterion of each evaluation instance, we employ a majority vote method. If at least two of the annotators share the same opinion, their preference is considered the final answer; otherwise, we consider the outputs of the two models to be tied. The results are shown in Figure 3.
Consistent with previous results, the AlpaCare-13B obtains better performance than PMC-13B in the human evaluation. The output generated by AlpaCare-13B is more accurate than that from PMC-13B, with 54% answers preferred by expert annotators. This result shows a superior medical capacity of AlpaCare. Beyond correctness, AlpaCare generates answers that are significantly more helpful compared to PMC-13B. 69% answers generated by AlpaCare are preferred by domain experts, highly exceeding the 11% preference rate for PMC-13B outputs, demonstrating better practical usability. The higher relative gain in helpfulness of AlpaCare compared to correctness is expected because the goal of IFT is to obtain the instruction-following ability of LLMs to align the model with various user needs, rather than gaining new knowledge (Zhou et al., 2023).
Figure 4: Analysis of diversity in the MedInstruct-52k. In panels (a-c), the top 20 entries for topic, view, and type are displayed, respectively. Panel (d) shows the distribution of instruction medical difficulty levels. Panel (e) analyzes linguistic diversity to depict the top 20 root verbs in the inner circle and their 4 primary direct noun objects in the outer circle in the generated instructions.
7 Analysis & Case Study
7.1 IFT Dataset Diversity Analysis
Training a model with diverse instructions enhances its ability to follow instructions (Wang et al., 2023b; Chen et al., 2023). However, current medical LLMs often have training data lacking in instructional diversity, typically using repetitive instructions across different instances (Li et al., 2023b; Han et al., 2023; Wu et al., 2023). To examine the diversity in our dataset, we plot the distributions of 4 key areas for instruction generation from MedInstruct-52k, shown in Figure 4 (a)- (d). Specifically, we present the top 20 topics, views and types and the difficulty levels from 1 to 5, offering insight into training data distribution. We further analyze instruction linguistic diversity by showing the root verbs and their corresponding direct-object nouns from each instruction. The top 20 root verbs and their 4 most common direct-object nouns are displayed in Figure 4 (e), representing 22% of the total dataset. Our findings show quite diverse medical intents and textual formats in our MedInstruct-52k.
Table 8: Quantitative comparison of linguistic diversity in medical instructional datasets. Comparing linguistic entropy of each IFT dataset for medical LLMs. The higher value represents better diversity.
ChatDoctor Medalpaca PMC Baize-H AlpaCare
Entropy 0 0 2.85 3.45 5.57
To quantitatively showcase our dataset’s diversity in comparison to IFT dataset of other medical LLMs, we calculate the linguistic entropy in the instructions of instruction-following datasets used for medical IFT models. Higher entropy values signify greater diversity. Specifically, we analyze the top 20 root verbs and their 4 primary direct noun objects for each dataset and calculate verb-noun pair entropy, as shown in Table 8. AlpaCare’s dataset, MedInstruct-52K, exhibits the highest entropy, underscoring its superior diversity, which in turn better elicits the model’s instruction-following capabilities in the IFT stage.
Figure 5: Case Study of 13B models of AlpaCare and PMC on (a) correctness and (b) helpfulness. Instruction key points and primary responses are highlighted in blue and yellow, respectively.
7.2 Generation Case Study
We randomly selected one win case from MedInstruct-test for correctness and another for helpfulness, as described in Section 6. Figure 5 displays the instructions and outputs of the base model, LLAMA-13B, and 13B medical models, AlpaCare and PMC.
Figure 5(a) showcases a case on correctness with a high medical difficulty level. LLAMA-13B struggles to provide correct answers to the questions in the instruction. In contrast, Medical LLMs demonstrate a marked improvement in the response accuracy. PMC offers a general overview, citing common causes like stress, hormonal imbalances, and nutritional deficiencies. However, it falls short by using ambiguous terms like "certain medical conditions" and "an underlying medical condition," which lack specificity and fail to deliver actionable medical insights. Conversely, AlpaCare provides a detailed analysis, pointing specific conditions such as hypothyroidism and iron deficiency anemia that could explain the patient’s symptoms and offering a nuanced understanding that significantly enhances the precision and utility of the guidance provided. Additionally, it emphasizes the importance of medical attention for severe symptoms, making its recommendations more actionable.
In the case of helpfulness illustrated in Figure 5(b), both the base model and medical models generate accurate medical information on Metoprolol from different perspectives. However, LLAMA and PMC do not adhere to the instruction to use a bullet-point format, which would enhance clarity. Conversely, AlpaCare clearly follows the instruction with well-structured formatting, demonstrating that fine-tuning with a diverse instruction-following dataset improves the ability to follow instructions and, in turn, increases helpfulness.
8 Conclusion
In conclusion, our study fills a gap in aligning LLMs with domain-specific user intents in the medical domain while managing costs. Specifically, we propose a semi-automated pipeline using GPT-4 and ChatGPT with a high-quality expert-curated seed set to create a diverse machine-generated medical IFT dataset, MedInstruct-52k, for LLM tuning. The resulting model, AlpaCare, demonstrates strong medical instruction-following and robust generalization abilities compared to previous medical LLMs. Our approach shows versatility across various metrics, including different LLM judges (such as ChatGPT and Claude-2), a range of base model families (like LLaMA-1 and LLaMA-2), and various model sizes (7B and 13B). This highlights the significant benefits of incorporating diverse data in medical AI model development.
9 Ethics Statement
This research demonstrates the potential of enhancing open-source LMs’ medical capacity and general applicability by distilling knowledge from more powerful "teacher" LLMs. We improve the smaller models’ ability to follow medical instructions and align with user intentions, offering potential benefits for future healthcare applications. Even though our model AlpaCare maintains comparable truthfulness scores compared to other baselines, but the rate of correct answers is still low, showing that our model is likely to hallucinate incorrect answers. Therefore, it is crucial to emphasize that this system is designed to serve as an assistant tool, complementing but not replacing the expertise and judgment of healthcare professionals. All outputs generated by this system must be rigorously validated by licensed medical practitioners before any practical application.
The transition of these LMs to practical medical scenarios — useful for healthcare professionals, patients, and other medical personnel — necessitates extensive research to ensure safety, privacy, and reliability. This development involves not only technical robustness but also a commitment to ethical standards. It includes comprehensive quality assessments in various clinical environments to ensure that the system meets the highest standards of accuracy and ethical conduct. Additionally, attention must be paid to the potential implications for patient privacy and data security, ensuring that all patient information is handled with the utmost confidentiality and in compliance with relevant data protection laws and healthcare regulations.
Acknowledgments
We gratefully acknowledge the generous financial support provided by the National Institutes for Health (NIH) grant NIH 7R01HL149670.
Appendix A Appendix
Limitations
Our approach utilizes ’teacher’ LLMs, such as GPT-4 and ChatGPT, to automatically generate medical instruction-response pair datasets, employing these teacher models as medical knowledge bases. However, this could result in hallucinations in the medical knowledge generation. To enhance the generation reliability, we aim to integrate LLMs with the internet and knowledge graphs in future work.
A.1 Medical Task Difficulty Level Scoring System
Table 9: Scoring system for evaluating the difficulty level of medical tasks.
Score Description
1 The fact is very basic. The answer becomes apparent immediately after reading the question, or it can be easily found through a direct internet search.
2 The fact is simple but may require a slight application of real-world knowledge, rephrasing, or extending the information to find the answer.
3 This involves facts that require more real-world application, dealing with practical and somewhat complicated situations. It may require more complex paraphrasing and/or communication skills, such as emotional support, psychological evaluations, and ethical considerations. The tested knowledge in this category can be quite challenging.
4 This level involves complicated medical facts. Answering questions at this level may require multi-step thinking processes. The questions might be lengthy and detailed, necessitating simplification for a clearer answer. This category might include most USMLE questions. It may also require a demonstration of enhanced emotional support, psychological evaluations, and ethical considerations. Questions might be based on vague symptom descriptions, making the diagnosis challenging, or involve recent advancements, publications, or current global health issues like pandemics.
5 This category involves complex medical knowledge applied to real-world, intricate situations. The questions are detailed and lengthy, often requiring simplification and multi-step thinking to answer. Some questions might be based on actual medical cases with challenging diagnoses and treatments. The symptom descriptions might be highly vague. Questions could also involve new technologies, recent publications, or current pandemics, requiring decision-making or choosing the best available option. Instructions might also necessitate the demonstration of humane care.
We introduce a clinician-crafted seed set to generate MedInstruct-52k and a free-form medical instruction evaluation set, MedInstruct-test . This set spans a medical difficulty scale ranging from 1 to 5, where 1 represents the easiest tasks and 5 indicates the most challenging ones. A clinician assessed the difficulty levels of all instances within both the seed set and MedInstruct-test based on the scoring system shown in Table 9 This system offers a refined dimension for prompting GPT-4 to produce tasks across varied difficulty levels and to evaluate medical proficiency of IFT models.
A.2 Prompt details for MedInstruct-52k generation
Here we provides prompts we use for query GPT-4 and ChatGPT for task and response generation.
Table 10: Task generation prompt
Your objective is to generate diverse medical-related tasks.
Here are the requirements:
1. Ensure that all tasks are related to the medical domain.
2. Craft tasks that encompass varied points of view, e.g. experts, students and patients, etc.
3. Maximize the range of task topics, e.g. diseases, treatment, diagnoses, epidemiology, pharmacology, pathophysiology, anatomy, genetics, medical education, etc.
4. Introduce different task formats, e.g. text generation, open Q&A, chat, rewrites, summarizations, classifications, USMLE style Q&A, multiple-choice Q&A, single-hop reasoning and multiple-hop reasoning etc.
5. All the formats specified in point 4 MUST be represented in the task you generate.
6. Create tasks with medical difficulty levels from 1 to 5, with 1 being the easiest and 5 the hardest.
7. Use diverse language in the instructions. For instance, combine questions with imperative forms.
8. Some instructions might require specific inputs. If an input is not necessary, such as with general instructions like "What are the side effects of COVID-19?", use "<noinput>" in the input field.
9. When provided, inputs must range between 50 to 200 words and offer detailed medical context , e.g. symptom descriptions, radiology reports, clinical notes, and exam questions, etc.
10. Generate a detailed and comprehensive input instead ask user-provided input.
11. Ensure USMLE style Q&A and multiple-choice Q&A tasks have both question and choices in input, and the question context should be detailed.
12. The USMLE-style question length must exceed 50 words.
13. Match instruction and input to the task’s perspective. Patient perspectives should be simple and in first person, while clinician views should have professional terminology.
14. Ensure the lengths of inputs for different tasks are notably distinct.
[…]
List of 15 tasks:
Seed task 1
Seed Task 2
Seed Task 3
Table 11: Output generation prompt
You are a medical expert tasked with answering various medical questions. You MUST generate your response based on the requirements.
Here are the requirements:
1. For multiple-choice, calculation, and classification problems, you can generate intermediate thinking steps if necessary; otherwise, provide the final answer directly.
2. All the intermediate thinking steps must be generated before final answer.
3. For multiple-choice questions, you MUST generate the answer choice in the following format: ‘The answer is (your choice).’ For example:
‘Choose the correct answer. Where in your body will you find the tibia bone? A) Arm B) Foot C) Skull D) Leg
The tibia bone is one of the two bones in the lower leg, the other being the fibula. The answer is D) Leg.’
4. For other types of questions, except multiple-choice, do not use the format mentioned in point 3.
task instruction
task input (if exist)
A.3 Training hyperparameter details
Table 12: AlpaCare hyperparameter setup.
Model Size Data Size GPUs Epoch LR Batch Size
7B 52k 4 40G A100 3 2e-5 128
13B 52k 4 80G A100 5 1e-5 128
A.4 More experimental results
A.4.1 General domain free-form instruction evaluation
Table 13: Comparison on general domain free-form instruction evaluation. A performance comparison between AlpaCare and IT baselines on AlpacaFarm on 4 distinct reference models: Text-davinci-003, GPT-3.5-turbo, GPT-4 and Claude-2. ‘AVG’ represents the mean performance score across all referenced models.
AlpacaFarm
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
Alpaca 38.7 20.6 14.5 16.9 22.7
ChatDoctor 37.4 20.3 13.1 14.0 21.2
Medalpaca 38.2 24.4 20.6 20.1 25.8
PMC 15.8 2.6 13.3 1.6 8.3
Baize-H 29.9 16.9 12.7 13.7 18.3
AlpaCare 56.4 38.6 34.2 33.7 40.7
A.4.2 Ablation study
Table 14: Medical free-from instruction evaluation results by using Claude-2 as judge.
iCliniq
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
Alpaca 40.7 33.4 18.5 14.2 26.7
ChatDoctor 24.5 23.8 11.3 9.9 17.4
Medalpaca 38.3 37.1 15.8 15.4 26.7
PMC 0 0 1.9 3.1 1.3
Baize-H 46.9 35.6 12.1 7.5 25.5
AlpaCare 64.5 46.8 26.9 17.1 38.8
MedInstruct
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
Alpaca 44.4 19.9 19.2 13.4 23.5
ChatDoctor 34.8 17.5 14.1 10.3 21.7
Medalpaca 41.4 19.3 18.0 12.5 23.1
PMC 2.8 1.5 0.1 0.7 1.8
Baize-H 40.1 25.1 21.8 15.2 19.8
AlpaCare 76.4 42.6 42.9 31.6 31.5
Table 15: Results on different LLM backbone across 4 reference models by using gpt-3.5-tubro as the judge. Comparing the performance of AlpaCare and Alpaca using different LLM backbones, with 4 distinct reference models.
iCliniq
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
LLaMA Alpaca 38.8 30.4 12.8 15.6 24.4
AlpaCare 66.6 50.6 47.4 49.7 53.6
LLaMA-2 Alpaca 45.8 36.3 18.2 20.8 30.3
AlpaCare 66.5 50.4 47.8 50 53.7
MedInstruct
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
LLaMA Alpaca 35.0 20.6 21.5 15.6 24.4
AlpaCare 67.6 48.8 47.4 49.7 53.5
LLaMA-2 Alpaca 39.6 22.7 26.4 18.5 26.8
AlpaCare 70.6 48.8 50.0 48.4 54.2
Table 16: Result comparison of 4 reference models on 13B instruction-tuned models.
iCliniq
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
Alpaca 46.7 37.0 19.6 21.7 31.3
Medalpaca 8.1 4.4 1.0 2.0 3.9
PMC 40.6 29.0 14.3 17.5 25.4
AlpaCare 66.7 51.2 48.2 50.2 54.4
MedInstruct
Text-davinci-003 GPT-3.5-turbo GPT-4 Claude-2 AVG
Alpaca 39.8 22.5 27.1 18.1 26.9
Medalpaca 0.2 0 0 0 0.1
PMC 44.9 31.9 32.8 29.2 34.7
AlpaCare 71.3 49.1 49.8 47.7 54.5